# KOALA++: Efficient Kalman-Based Optimization of Neural Networks with Gradient-Covariance Products

### v1

@@ Acknowledgements
1
Appendix A KOALA
In this section, we are going to revisit the original KOALA framework, which was only briefly introduced in the main text. KOALA [3] is a Kalman-inspired optimization algorithm that interprets gradient-based updates as a recursive state estimation process. To ensure scalability in high-dimensional deep learning settings, the original work proposed two tractable variants: KOALA-V and KOALA-M. These variants differ in how they approximate the parameter covariance and in whether they incorporate momentum dynamics. In what follows, we provide a self-contained and detailed exposition of both KOALA-V and KOALA-M, including their derivations, update rules, and practical considerations. This not only offers a clearer understanding of KOALA as a foundational method, but also sets the stage for the improvements introduced in KOALA++.
A.1 Risk Minimization as Loss Adaptivity
The core idea behind KOALA is to cast the optimization of deep neural networks as a state estimation problem under uncertainty. Specifically, the training of neural networks is considered as minimizing a noisy risk function ${L}_{k}$, which is a stochastic estimate of the empirical risk ${\cal L}$. Motivated by the central limit theorem, KOALA assumes that the mini-batch risk ${L}_{k}$ can be modeled as a Gaussian random variable with scalar noise:
@@ A.1 Risk Minimization as Loss Adaptivity
where $\theta$ denotes the model parameters and $R_{k}$ is the variance of the observation noise.
Rather than optimizing $\min_{\theta}{\cal L}(\theta)$ directly, KOALA reformulates the problem as adapting the model parameters such that the stochastic risk ${L}_{k}(\theta_{k})$ matches a desired target risk ${L}^{\text{target}}_{k}$, i.e.,
@@
This formulation enables the use of Kalman filtering, which provides a principled way to recursively update the estimate of $\theta_{k}$ by fusing prior knowledge with new, noisy observations. The underlying system can be described as a standard dynamical model:
@@
${L}^{\text{target}}_{k}$ $={L}_{k}(\theta_{k})+v_{k},\quad v_{k}\sim\mathcal{N}(0,R_{k})$ (24)
@@ A.2 KOALA-V (Vanilla)
The idea behind KOALA-V is remarkably natural, which is directly appling the standard Kalman filtering update equations to the system in Equations (23) and (24). However, directly computing the posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$ is infeasible for high-dimensional models. To address this, KOALA-V introduces the following approximations:
@@ A.3 KOALA-M (Incorporating Momentum Dynamics)
$p_{k}$ $=\kappa\,p_{k-1}+u_{k-1},\quad u_{k-1}\sim\mathcal{N}(0,U),$ (31)
@@ Appendix B Mathematical Foundations of KOALA++
In the main paper, we provided both the non-symmetric and symmetric solution for $P_{k-2}$. In this section, we provide the detailed derivations of those. For notational simplicity, where possible we drop the subscript in $k$, and we can restate the optimization problem as follows:
@@ B.1 Vanilla Optimal Solution of Covariance Matrix P
To solve this constrained optimization problem, we introduce a Lagrange multiplier $\lambda\in\mathbb{R}^{1\times n}$ and construct the corresponding Lagrangian function:
$\mathcal{L}(P,\lambda)=\|P\|_{F}^{2}+\lambda(HP-v)^{\top}.$ (33)
@@
$\frac{\partial\mathcal{L}(P,\lambda)}{\partial P}=2P+H^{\top}\lambda=0\quad\Rightarrow\quad P=-\frac{1}{2}H^{\top}\lambda.$ (34)
@@
$H\left(-\frac{1}{2}H^{\top}\lambda\right)$ $=v\Rightarrow\lambda=-2\frac{v}{HH^{\top}}.$ (35)
Thus, the optimal solution for $P$ is obtained by substituting back the expression for $\lambda$:
@@
This solution is referred to as the vanilla optimal solution (asymmetric) because it does not impose symmetry constraints explicitly on $P$. In the next subsection, we derive a symmetric solution variant that respects the symmetry requirement typically imposed on covariance matrices.
@@ B.2 Symmetric Optimal Solution of Covariance Matrix P
For derivation of the symmetric solution, we use the same Lagrangian framework introduced earlier, but now explicitly incorporate the symmetry constraint into the formulation. The Lagrangian function, considering the symmetry constraint, remains the same as before (see Equation (33)).
@@
$\frac{\partial\mathcal{L}}{\partial P}(\text{Symmetric})=\left[\frac{\partial\mathcal{L}}{\partial P}\right]+\left[\frac{\partial\mathcal{L}}{\partial P}\right]^{T}-\text{diag}\left[\frac{\partial\mathcal{L}}{\partial P}\right]=0$ (37)
@@
$02P+\lambda^{T}H+2P^{T}+H^{T}\lambda-\text{diag}(2P+\lambda^{T}H)=$ 0 (38)
Since $P$ is symmetric, we can rewrite the Equation (38) as:
$4P+\lambda^{T}H+H^{T}\lambda-\text{diag}(2P+\lambda^{T}H)=0,$ (39)
which gives the equations as follows:
$\begin{cases}4P+\lambda^{\top}H+H^{\top}\lambda=0&\text{(off diagonal)}\\ 2P+H^{\top}\lambda=0&\text{(on diagonal)}\end{cases}$ (40)
We observe that the off-diagonal solution satisfies the on-diagonal equation on the diagonal. Therefore, we only need to solve the off-diagonal case. From the off-diagonal constraint $04P+\lambda^{\top}H+H^{\top}\lambda=$ 0, we have $P=-\frac{1}{4}\left(\lambda^{\top}H+H^{\top}\lambda\right)$. By substituting in $HP=v$, we obtain
$HP=-\frac{1}{4}(H\lambda^{\top}H+HH^{\top}\lambda)=v.$ (41)
We can solve $\lambda$ as
$\lambda=-4v(H^{\top}H+HH^{\top}I)^{-1},$ (42)
@@
By letting $A=HH^{\top}I$, $U=H^{\top}$, and $V=H$, we find that:
$(HH^{\top}I+H^{\top}H)^{-1}=\frac{1}{HH^{\top}}\left(I-H^{\top}\cdot\frac{1}{2}H\cdot\frac{1}{HH^{\top}}\right)$ (44)
This provides an expression for $\lambda$ without matrix inverses, i.e.,
$\lambda=-4v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right).$ (45)
@@ B.3 Eigenvalues Analysis of the Optimal Matrix $P^{*}$
$P^{*}=\frac{1}{HH^{T}}(H^{\top}v+v^{T}H)-\frac{Hv^{T}}{(HH^{T})^{2}}H^{T}H.$ (47)
To analyze the definiteness of $P^{*}$, we consider its eigenvalues. Let $W\in\mathbb{R}^{n\times 1}$ be an eigenvector of $P^{*}$, satisfying
$\lambda W=P^{*}W$ (48)
@@
$P^{*}W$ $=\frac{1}{A}(H^{T}v\alpha H^{T}+H^{T}v\beta v^{T}+v^{T}H\alpha H^{T}+v^{T}H\beta v^{T})$ (52)
$-\frac{1}{A^{2}}(Hv^{T}H^{T}H\alpha H^{T}+Hv^{T}H^{T}H\beta v^{T})$ (53)
$=\frac{1}{A}(\alpha BH^{T}+\beta CH^{T}+\alpha Av^{T}+\beta Bv^{T})-\frac{1}{A^{2}}\alpha BAH^{T}-\frac{1}{A^{2}}\beta B^{2}H^{T}$ (54)
$=\frac{\beta C}{A}H^{T}+\alpha v^{T}+\frac{\beta B}{A}v^{T}-\frac{\beta B^{2}}{A^{2}}H^{T}$ (55)
$=\frac{\beta(AC-B^{2})}{A^{2}}H^{T}+\frac{\alpha A+\beta B}{A}v^{T}.$ (56)
@@
$\alpha\lambda$ $=\dfrac{\beta(AC-B^{2})}{A^{2}}$ (57)
$\beta\lambda$ $=\dfrac{\alpha A+\beta B}{A}.$ (58)
If $\alpha=0$, then from Equation (57) we know that $\beta=0$ or $AC-B^{2}=0$. However, $\beta=0$ implies that $W=\alpha H^{\top}+\beta v^{\top}=0$, which contradicts the assumption that $W$ is a non-zero eigenvector. On the other hand, by the Cauchy-Schwarz inequality, $AC-B^{2}=0$ implies that the two directions $H^{\top}$ and $v^{\top}$ become linearly dependent. In this case, the eigenvalues of $P^{*}$ are 00 and $a$, where $H^{\top}=av^{\top}$, due to the linear dependence. The linear dependence is, however, very unlikely in practice, as, in general $H$ and $v$ span a two-dimensional subspace.
Therefore, we must have $\alpha\neq 0$. In this case, the eigenvalue $\lambda$ must satisfy
$\lambda=\frac{\beta}{\alpha}\cdot\frac{AC-B^{2}}{A^{2}}.$ (59)
Now, we need to solve the linear system to examine the sign of $\frac{\beta}{\alpha}$
$(AC-B^{2})\left(\frac{\beta}{\alpha}\right)^{2}-AB\frac{\beta}{\alpha}-A^{2}=0.$ (60)
@@
$\frac{\beta}{\alpha}=\frac{AB\pm\sqrt{A^{2}B^{2}+4(AC-B^{2})A^{2}}}{2(AC-B^{2})}.$ (61)
@@
$\lambda_{1,2}=\frac{AB\pm\sqrt{A^{2}B^{2}+4(AC-B^{2})A^{2}}}{2A^{2}}.$ (62)
@@
Appendix C Ablations
C.1 Effect of Batch Size on KOALA++ Performance

### v2

@@ 80. Guidelines:
Appendix A KOALA
In this section, we are going to revisit the original KOALA framework, which was only briefly introduced in the main text. KOALA [4] is a Kalman-inspired optimization algorithm that interprets gradient-based updates as a recursive state estimation process. To ensure scalability in high-dimensional deep learning settings, the original work proposed two tractable variants: KOALA-V and KOALA-M. These variants differ in how they approximate the parameter covariance and in whether they incorporate momentum dynamics. In what follows, we provide a detailed and self-contained exposition of both KOALA-V and KOALA-M, including their derivations, update rules, and practical considerations. This not only offers a clearer understanding of KOALA as a foundational method, but also sets the stage for the improvements introduced in KOALA++.
A.1 Risk Minimization as Loss Adaptivity
@@
$\theta$ denotes the model parameters and $R_{k}$ is the variance of the observation noise.
Rather than optimizing $\min_{\theta}{\cal L}(\theta)$ directly, KOALA reformulates the problem by adapting the model parameters so that the stochastic risk ${L}_{k}(\theta_{k})$ matches a desired target risk ${L}^{\text{target}}_{k}$, i.e.,
@@
This formulation enables the use of Kalman filtering, which provides a principled way to recursively update the estimate of $\theta_{k}$ by combining prior knowledge with new and noisy observations. The underlying system can be described as a standard dynamical model:
@@
${L}^{\text{target}}_{k}$ $={L}_{k}(\theta_{k})+v_{k},\quad v_{k}\sim\mathcal{N}(0,R_{k}).$ (24)
@@ A.2 KOALA-V (Vanilla)
The idea behind KOALA-V is remarkably natural, which is to directly apply the standard Kalman filtering update equations to the system in Equations (23) and (24). However, directly computing the posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$ is infeasible for high-dimensional models. To address this, KOALA-V introduces the following approximations:
@@ A.3 KOALA-M (Incorporating Momentum Dynamics)
$p_{k}$ $=\kappa\,p_{k-1}+u_{k-1},\quad\quad\quad\quad u_{k-1}\sim\mathcal{N}(0,U),$ (31)
@@ Appendix B Mathematical Foundations of KOALA++
In the main paper, we provided both the non-symmetric and symmetric solutions for $P_{k-2}$. In this section, we provide the detailed derivations of those. For notational simplicity, where possible, we drop the subscripts in $k$, and we can restate the optimization problem as follows:
@@ B.1 Vanilla Optimal Solution of Covariance Matrix P
To solve this constrained optimization problem, we introduce a Lagrange multiplier $\mu\in\mathbb{R}^{1\times n}$ and construct the corresponding Lagrangian function:
$\mathcal{L}(P,\mu)=\|P\|_{F}^{2}+\mu(HP-v)^{\top}.$ (33)
@@
$\frac{\partial\mathcal{L}(P,\mu)}{\partial P}=2P+H^{\top}\mu=0\quad\Rightarrow\quad P=-\frac{1}{2}H^{\top}\mu.$ (34)
@@
$H\left(-\frac{1}{2}H^{\top}\mu\right)$ $=v\Rightarrow\mu=-2\frac{v}{HH^{\top}}.$ (35)
Thus, the optimal solution for $P$ is obtained by substituting back the expression for $\mu$:
@@
This solution is referred to as the vanilla optimal solution (asymmetric) because it does not explicitly impose symmetry constraints on $P$. In the next subsection, we derive a symmetric solution variant that respects the symmetry requirement typically imposed on covariance matrices.
@@ B.2 Symmetric Optimal Solution of Covariance Matrix P
For derivation of the symmetric solution, we use the same Lagrangian framework, but now explicitly incorporate the symmetry constraint into the formulation. The Lagrangian function, considering the symmetry constraint, remains the same as before (see Equation (33)).
@@
$\frac{\partial\mathcal{L}}{\partial P}(\text{Symmetric})=\left[\frac{\partial\mathcal{L}}{\partial P}\right]+\left[\frac{\partial\mathcal{L}}{\partial P}\right]^{\top}-\text{diag}\left[\frac{\partial\mathcal{L}}{\partial P}\right]=0$ (37)
@@
$2P+\mu^{\top}H+2P^{\top}+H^{\top}\mu-\text{diag}(2P+\mu^{\top}H)=0$ (38)
Since $P$ is symmetric, we can rewrite Equation (38) as:
$4P+\mu^{\top}H+H^{\top}\mu-\text{diag}(2P+\mu^{\top}H)=0,$ (39)
which gives the following equations
$\begin{cases}4P+\mu^{\top}H+H^{\top}\mu=0&\text{(off diagonal)}\\ 2P+H^{\top}\mu=0&\text{(on diagonal).}\end{cases}$ (40)
We observe that the off-diagonal solution satisfies the on-diagonal equation on the diagonal. Therefore, we only need to solve the off-diagonal case. From the off-diagonal constraint $4P+\mu^{\top}H+H^{\top}\mu=0$, we have $P=-\frac{1}{4}\left(\mu^{\top}H+H^{\top}\mu\right)$. By substituting in $HP=v$, we obtain
$HP=-\frac{1}{4}(H\mu^{\top}H+HH^{\top}\mu)=v.$ (41)
We can solve $\mu$ as
$\mu=-4v(H^{\top}H+HH^{\top}I)^{-1},$ (42)
@@
By letting $A=HH^{\top}I$, $U=H^{\top}$, and $V=H$, we find the following expression
$(HH^{\top}I+H^{\top}H)^{-1}=\frac{1}{HH^{\top}}\left(I-H^{\top}\cdot\frac{1}{2}H\cdot\frac{1}{HH^{\top}}\right).$ (44)
This provides an expression for $\mu$ without matrix inverses, i.e.,
$\mu=-4v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right).$ (45)
@@ B.3 Eigenvalues Analysis of the Optimal Matrix $P^{*}$
$P^{*}=\frac{1}{HH^{\top}}(H^{\top}v+v^{\top}H)-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H.$ (47)
To analyze the definiteness of $P^{*}$, we consider its eigenvalues. Let $w\in\mathbb{R}^{n\times 1}$ be an eigenvector of $P^{*}$ and $\lambda$ its corresponding eigenvector, then,
$\lambda w=P^{*}w.$ (48)
@@
$P^{*}w$ $=\frac{1}{x}(H^{\top}v\alpha H^{\top}+H^{\top}v\beta v^{\top}+v^{\top}H\alpha H^{\top}+v^{\top}H\beta v^{\top})$ (52)
$-\frac{1}{x^{2}}(Hv^{\top}H^{\top}H\alpha H^{\top}+Hv^{\top}H^{\top}H\beta v^{\top})$ (53)
$=\frac{1}{x}(\alpha yH^{\top}+\beta zH^{\top}+\alpha xv^{\top}+\beta yv^{\top})-\frac{1}{x^{2}}\alpha yxH^{\top}-\frac{1}{x^{2}}\beta y^{2}H^{\top}$ (54)
$=\frac{\beta z}{x}H^{\top}+\alpha v^{\top}+\frac{\beta y}{x}v^{\top}-\frac{\beta y^{2}}{x^{2}}H^{\top}$ (55)
$=\frac{\beta(xz-y^{2})}{x^{2}}H^{\top}+\frac{\alpha x+\beta y}{x}v^{\top}.$ (56)
@@
$\alpha\lambda$ $=\dfrac{\beta(xz-y^{2})}{x^{2}}$ (57)
$\beta\lambda$ $=\dfrac{\alpha x+\beta y}{x}.$ (58)
If $\alpha=0$, then from Equation (57) we know that $\beta=0$ or $xz-y^{2}=0$. However, $\beta=0$ implies that $w=\alpha H^{\top}+\beta v^{\top}=0$, which contradicts the assumption that $w$ is a non-zero eigenvector. On the other hand, by the Cauchy-Schwarz inequality, $xz-y^{2}=0$ implies that the two directions $H^{\top}$ and $v^{\top}$ become linearly dependent. In this case, the eigenvalues of $P^{*}$ are 0 and $a$, where $H^{\top}=av^{\top}$, due to the linear dependence. The linear dependence is, however, very unlikely in practice, as, in general $H$ and $v$ span a two-dimensional subspace.
Therefore, we very likely have $\alpha\neq 0$. In this case, the eigenvalue $\lambda$ must satisfy
$\lambda=\frac{\beta}{\alpha}\cdot\frac{xz-y^{2}}{x^{2}}.$ (59)
Now, we need to solve the following equation to examine the sign of $\frac{\beta}{\alpha}$
$(xz-y^{2})\left(\frac{\beta}{\alpha}\right)^{2}-xy\frac{\beta}{\alpha}-x^{2}=0.$ (60)
@@
$\frac{\beta}{\alpha}=\frac{xy\pm\sqrt{x^{2}y^{2}+4(xz-y^{2})x^{2}}}{2(xz-y^{2})}.$ (61)
@@
$\lambda_{1,2}=\frac{xy\pm\sqrt{x^{2}y^{2}+4(xz-y^{2})x^{2}}}{2x^{2}}.$ (62)
@@
Appendix C Convergence Behavior of KOALA++
C.1 On the Stability and Implicit Positive Semi-Definiteness of KOALA++
While the current formulation of KOALA++ does not explicitly enforce the positive semi-definite (PSD) constraint on the covariance-like matrix, we empirically observe that the optimizer remains numerically stable and convergent across diverse architectures and datasets. This suggests that, in practice, the lack of an explicit PSD guarantee does not adversely affect convergence behavior.
This is probably because, when we are calculating the eigenvalues in Equation (62), the quantities $\lambda_{1,2}$ are obtained from the least-squares solution of the auxiliary matrix $P_{k-2}$, which represents the optimal estimate under the minimal-norm fitting objective. In other words, the eigenvalues in Equation (62) correspond to an analytically fitted surrogate of the covariance update, rather than the actual covariance-like matrix used in the algorithmic recursion. In practice, the true update matrix—hereafter referred to as the ground truth matrix—is defined by integrating the solution from Equation (12) into (9), then we have:
$v_{k}$ $=H_{k}\left(\frac{H^{\top}v+v^{\top}H}{\|H\|^{2}}-\frac{(Hv^{\top})(H^{\top}H)}{\|H\|^{4}}\right.\left.+QI-\frac{(v^{\top}+QH^{\top})(v+QH)}{H(v^{\top}+QH^{\top})+R}\right).$ (63)
We call the ground-truth matrix $M_{k}$ and express it as follows
$M_{k}$ $=\frac{H^{\top}v+v^{\top}H}{\|H\|^{2}}-\frac{(Hv^{\top})(H^{\top}H)}{\|H\|^{4}}+QI-\frac{(v^{\top}+QH^{\top})(v+QH)}{H(v^{\top}+QH^{\top})+R}.$ (64)
Based on this formulation, we can further examine the directional definiteness of the update. Specifically, by evaluating the product $H_{k}v_{k}^{\top}$, we can infer whether the update direction induced by $M_{k}$ corresponds to a positive-definite step along the current gradient direction. In particular, when
$H_{k}v_{k}^{\top}=H_{k}M_{k}H_{k}^{\top}>0,$ (65)
the update matrix $M_{k}$ can be considered positive definite with respect to the local curvature defined by $H_{k}$. This provides a practical criterion for assessing the stability and consistency of the covariance update in KOALA++.
Empirically, we verify this property on three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet32. For each case, we monitor the angle between the curvature vector $H_{k}$ and the update direction $v_{k}$ throughout the training. As shown in Figure 4, the angle remains consistently acute ($90^{\circ}$) in all training stages, which implies that $H_{k}v_{k}^{\top}>0$, thus confirming the numerical positive semi-definiteness of $M_{k}$ and the empirical stability of the update rule.
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet32
Figure 4: Empirical verification of the positive semi-definiteness of $M_{k}$ in KOALA++. The plots show the evolution of the angle between the curvature vector $H_{k}$ and the update direction $v_{k}$ across training for three datasets. The consistently acute angles ($90^{\circ}$) indicate that $H_{k}$ and $v_{k}$ remain directionally aligned, confirming the effective PSD behavior of $M_{k}$ in practice.
However, from a theoretical standpoint, ensuring positive-semidefiniteness could further strengthen the stability guarantees of the method. In particular, one promising direction is to modify the optimization objective so that the covariance estimate minimizes its distance—for example, in the Frobenius norm—to a structured PSD matrix such as a scaled identity. Such a formulation naturally leads to a convex optimization problem with a unique and bounded solution, thereby eliminating ambiguity and improving theoretical soundness.
Although this PSD-enforced variant is beyond the current scope of our study, we view it as an important extension to future work aimed at establishing stronger convergence guarantees for KOALA++.
C.2 Convergence Characteristics of KOALA++ vs. Quasi-Second-Order Optimizers
To better understand how KOALA++ differs from quasi-second-order optimizers in convergence behavior, we conduct a controlled comparison on CIFAR-10 using a ResNet-50 backbone trained for 100 epochs under a cosine learning rate scheduler.
Although the Shampoo optimizer [8] is a well-established quasi-second-order method, its computational overhead makes direct comparison impractical under identical epoch budgets. To ensure a fair and efficient evaluation, we instead adopt Muon [14], a more recent and scalable curvature-aware optimizer that follows a similar quasi-second-order philosophy but operates with significantly lower wall-clock cost. This substitution allows us to better examine the convergence profile of KOALA++ against comparable quasi-second-order approaches.
(a) Log-scale Test Loss vs Epochs
(b) Top-1 Validation Error vs Epochs
Figure 5: CIFAR-10 with ResNet18 (Left) and ResNet50 (Right) trained for 100 epochs under a cosine learning-rate scheduler. Validation Top-1 errors (the inset highlights the late-epoch region) are reported.
As shown in Figure 5, Muon achieves faster convergence in the early phases, while KOALA++ initially exhibits larger fluctuations but gradually stabilizes and reaches comparable or better final accuracy. We attribute this difference to the distinct curvature modeling mechanisms of the two optimizers. Muon performs explicit matrix preconditioning with guaranteed positive semi-definite (PSD) curvature estimates, whereas KOALA++ relies on an implicit least-squares projection derived from scalar observations. Since the resulting covariance-like matrix is not explicitly constrained to be PSD, KOALA++ may require several epochs to implicitly correct unstable directions before converging to a well-conditioned update regime. This interpretation remains a hypothesis, and a rigorous theoretical analysis of the PSD property and its impact on convergence stability is left to future work.
Appendix D Ablations
D.1 Effect of Batch Size on KOALA++ Performance